# Application of Machine Learning in Agriculture: Recent Trends and Future Research Avenues

### v1

@@ 5.4 Machine Learning Application in Disease Detection
34 Classifying tomato ripeness in the design of a tomato harvesting robot 2018 Improved deep learning-based classification method that improves the accuracy and scalability of tomato ripeness with a small amount of training data. CNN,DNN Less parameter calculation and higher accuracy. over-fitting probelm [52]
@@
35 Automated fruit harvesting robot 2019 This research introduces an approach for the identification and harvesting of fruits using a robotic arm. CNN, R-CNN, VGG, YOLO The algorithm for fruit position detection and harvesting can be automated High cost of model [96]
@@
36 RTFD real time fruit detection 2019 This study employs novel convolutional deep learning techniques based on single-shot detectors for fruit (apple and pear) detection and counting within the tree canopy. CNN, YOLO Redues fruit losses The limitation of the model in differentiating between two objects of the same category in the same grid cell may result in inaccurate fruit counting in some cases [97]
@@
37 Method for mango detection and counting 2019 This study addresses the challenge of stable fruit detection in varying illumination conditions present in open farm environments. CNN MangoNet, a deep CNN architecture using sematic segmentation Stable fruit detection in varying illumination conditions can improve efficiency and accuracy of fruit counting and yield estimation in open farm environments This work only focused on one fruit, so the model will not perform well for other fruits. [53]
@@
38 Fruit feature detection 2019 Use of (DCNN) to optimize features for classification and segmentation tasks in agricultural navigation VGG16 (OxfordNet),Dcnn The use of DCNNs and VGG16 for feature optimization and classification improve accuracy in agricultural navigation tasks. Models works for specific crops only [98]
@@
39 Harvesting sweet pepper fruit in greenhouses 2019 The framework presented in this study incorporates three classification models that enable real-time classification of date fruit images based on their type, maturity, and decision regarding harvesting. AlexNet and VGG-16,CNN The model achieves best accuracy in fuits classiying The fine-tuning process of the VGG-16 model require significant computational resources and expertise to achieve optimal results. [54]
@@
40 Fruit Yield 2019 This paper presents the deep multifaced system to increase banana fruits yield LSTM, DL The model achieved better performance Acquiring and curating such data can be time-consuming and resource-intensive. [99]
@@
41 Fruit Detection 2020 FD should be robust to varying outdoor conditions with high signal to noise ratio. SD should help in estimating the angle at which the fruit is oriented around the stem for effective severing. Semantic Segmentation using FCN, Adaptive Image thresholding using video images using reinforced learning with decaying e-greedy algorithm FD system’s robustness and ability to handle varying outdoor conditions contribute to accurate fruit detection Computational complexity [100]
@@
42 Fruit picking robots, 2020 The system possesses the ability to train with multiple fruit types, enabling accurate classification based on factors such as size and ripeness. Improved efficiency Multiple fruit crops support Efficiency can be improved [55]
@@
43 Fruit picking 2020 This article presents a dual-arm aubergine harvesting robot consisting of two robotic arms configured in an anthropomorphic manner to optimize the dual workspace. Cubic SVM, water shed transformation, Point cloud extraction The use of a dual-arm configuration in the aubergine harvesting robot allows for optimized workspace Computational demands may limit real-time performance, requiring specialized hardware. [101]
@@
44 Fruit Detectetion 2020 The system demonstrates effectiveness in detecting fruits even in environments where leaves or branches have similar colors, accommodating variability in fruit shape and irregularities in the growing environment. CNN, SVM Improved cucumber harvesting efficiency through effective fruit detection and high recognition accuracy. limited adaptability for diverse agriculture applications [102]
@@
45 Fruit Detectetion 2020 To locate & harvest cotton balls in field conditions. Deep learning with Tiny Yolov3 with 7 conv. layers & CLoDSA based image augmentation Over $77\%$ Action Success Ratio ASR and over $97\%$ Detection Performance Future design and development research should also include alternative energy sources to decrease energy costs. [103]
@@
46 Fruit picking 2020 The paper introduces a harvesting robot designed specifically for plucking tea leaves, utilizing Jaco 2 robotic arm Jaco 2 based leaf plucking using ProMP Increased efficiency and reduced labor in tea production. Found limitations in handling branches with varying stiffness. [104]
@@
47 Autonomous path detection 2020 The paper presents a developing a system for autonomous path detection in semi-structured environments to enhance navigation and operation efficiency. CNN Potential for cost savings by minimizing human labor and errors. Poor lighting, occlusions, and complex surroundings can limit the accuracy of autonomous path detection. [105]
@@
48 Fruit Detection 2020 Improved detection of small harvesting cups using gray scale-depth and color images, with practical lighting-independent applications. Faster RCNN with MobileNetV2 Accurate detection of small harvesting cups in a wide range of lighting conditions, enabling more efficient and effective operations in diverse agricultural environments. Dependency on color images for tapping position detection. [106]
@@
49 Fruit Detection 2023 In this research, we propose lightweight fruit detection algorithm for edge CPUs, enabling 19 FPS detection speed in RTFD-CPU smartphone app.. DNN,PicoDet-S model The proposed RTFD has great potential for intelligent picking machines, Sensitivity to environmental factors that may affect fruit detection accuracy. [107]
@@
50 Post harvesting disease 2011 The system automatically extracts quality features from external data using image processing. CNN saves time and labor cost It has both hardware and software, space taking [56]
51 Post harvesting disease 2017 The system incorporates a prediction system for warehouse managers, recommending optimal dispatch sequences and transportation conditions for minimal storage and transportation loss. IoT,ML Optimized stock dispatch, reduced loss, and real-time monitoring empower farmers for efficient crop transportation and management. Security concerns [57]
52 Post harvesting disease 2019 Image processing technique used to classify mangoes based on their condition, considering three different phases. ML; DOF; Machine Vision; Articulated Arm; Fuzzy Logic Edge Detection; RGB Eliminating the need for manual labor while preserving the fruits features and appearance. The model only takes strong features to calculate and predicts the output of the data. [58]
53 Post harvesting disease 2019 The proposed system achieved detecting calyx and stalk scars in both defected and healthy tomatoes. RBF SVM Highest accuracy The model is only for tommatoes [59]
@@
54 Better post harvest handling 2020 Improving harvest quality and productivity for farmers through innovative solutions. Computer vision DL, YOLOV3 Assisting in the identification of healthy apples by distinguishing them from those with defects. Limitation in triaing diffrent type of data [108]
@@
55 Better post harvest handling 2020 A ML approach utilizes radar backscattering data to accurately estimate grain quantity in silos. d KNN Easy to apply in diffrerent grains Challenges in accuaring correct radar data [60]
@@
56 Food secuiritry 2021 A hierarchical model is proposed, integrating ontology-enabled IoT, to distinguish and separate healthy Sekai-ichi apples using Boosted Continuous Non-spatial Whole Attribute Extraction (BCNAE). IoT,ML Achieving higher accuracy and throughput while minimizing processing time. Computational complexity and accuracy-speed trade-off are potential drawbacks when aiming for higher accuracy, throughput, and speed. [109]
@@
57 Better post harvest handling 2022 The review explores the use of sensors, IoT principles, and ML technologies in post-harvest grain monitoring, identifying potential advancements and resource utilization. IoT Enhanced grain quality monitoring Collecting sensor data requires robust measures to protect sensitive information. [61]
58 Food security 2022 The paper has investigated household-level food security RF and CHAID decision tree offers an efficient and reliable data analysis method. Prone to over-fitting The effectiveness of the model vary across different locations. [110]
@@
59 Food security 2022 The paper has proposed detection and prediction models for identifying post-harvest deterioration zones in stored apple fruits. CNN-based U-Net, Deep-Lab , and Mask R-CNN The models enable retailers to maintain better quality control of their apple fruits, ensuring customer satisfaction. The models specifically designed for stored apple fruits only [111]
@@
60 Food secuiritry 2022 The paper has presented Drone technology for food supply chain management . DL,ML Reduces costs associated with traditional packing and delivery methods. Drons have limited payload capcity,battery life issues [112]
61 Post harvesting disease 2023 The paper introduces a knowledge-based expert system utilizing Bayesian network for diagnosing post-harvest diseases in apple. Bayesian network Enhance the trust in the system and improve user understanding. Implementing the application for real-world field e [113]
@@
Table 6 presents an overview of recent applications of machine learning techniques in the realm of agricultural practices, with a specific emphasis on genetic changes and plant breeding. The table chronicles research spanning from 2018 to 2023 and encapsulates various machine learning models. These models have been leveraged to accelerate crop improvement, aid in precision decision-making for speed breeding, enhance climate-resilient crop breeding, and even for automated plant phenotyping. Several references, such as [114] and [115], further underscore the significance of machine learning in advancing modern agricultural practices and outcomes.
@@
Plant Breeding 2019 Nueral Network, Explainable AI The author has presented state of genomics and field phenomics, and proposes a workable path to improvement. [114]
Plant Breeding 2023 Ensemble learning, Big data, Random forest ML has used in analysing genetic data in plant breeding, accerelating the crop improvemment, developing new varities and food data sequirity. [115]
@@
Integrating Speed Breeding 2022 Deep learning, Neural Network The author has utilized ML in speeding plant breeding and enabling precision decesion making. [116]
@@
Climate-Resilient Smart-Crop Breeding 2022 Machine learning The author has reviewed machine learning application in crop breeding. [117]
Plant Breeding 2020 Nueral Network, SVM, RF Discuused about improving new layers of plant breeding [118]
@@
Plant Breeding 2021 multilayer perceptron, SVM, and RF The ML models have been used for improving crop yield of soyabean [119]
@@
Plant Breeding 2019 RF Predict in-season seed yield (SY) in soybean based on high-dimensional phenotypic trait data [120]
@@
Plant Breeding 2019 SVM, DL The paper compares the performance of MLP and SVM methods to the Bayesian threshold genomic best linear unbiased prediction (TGBLUP) model for genomic selection in plant breeding. [121]
@@
Estimating Phenotypic Traits 2018 Regression, DL Using color images and elevation maps, machine learning is employed to estimate emergence and biomass traits in wheat field plots. [122]

### v2

@@ 3.3 Machine Learning Applications in Post-Harvesting
28 Classifying tomato ripeness in the design of a tomato harvesting robot 2018 Improved deep learning-based classification method that improves the accuracy and scalability of tomato ripeness with a small amount of training data. $\bullet$ CNN
$\bullet$ DNN
$\bullet$ Less parameter calculation and higher accuracy $\bullet$ Over-fitting problem [64]
@@
29 Automated fruit harvesting robot 2019 This research introduces an approach for the identification and harvesting of fruits using a robotic arm. $\bullet$ CNN
$\bullet$ R-CNN
$\bullet$ VGG
$\bullet$ YOLO
$\bullet$ The algorithm for fruit position detection and harvesting can be automated $\bullet$ High cost of model [167]
@@
30 RTFD real time fruit detection 2019 This study employs novel convolutional deep learning techniques based on single-shot detectors for fruit (apple and pear) detection and counting within the tree canopy. $\bullet$ CNN
$\bullet$ YOLO
$\bullet$ Redues fruit losses $\bullet$ The limitation of the model in differentiating between two objects of the same category in the same grid cell may result in inaccurate fruit counting in some cases [149]
@@
31 Method for mango detection and counting 2019 This study addresses the challenge of stable fruit detection in varying illumination conditions present in open farm environments. $\bullet$ CNN $\bullet$ MangoNet, a deep CNN architecture using semantic segmentation Stable fruit detection in varying illumination conditions can improve efficiency and accuracy of fruit counting and yield estimation in open farm environments $\bullet$ This work only focused on one fruit, so the model will not perform well for other fruits. [61]
@@
32 Fruit feature detection 2019 Use of (DCNN) to optimize features for classification and segmentation tasks in agricultural navigation. $\bullet$ VGG16 (OxfordNet)
$\bullet$ Dcnn
$\bullet$ The use of DCNNs and VGG16 for feature optimization and classification improve accuracy in agricultural navigation tasks. $\bullet$ Models work for specific crops only [63]
@@
33 Harvesting sweet pepper fruit in greenhouses 2019 The framework presented in this study incorporates three classification models that enable real-time classification of date fruit images based on their type, maturity, and decision regarding harvesting. $\bullet$ AlexNet and VGG-16
$\bullet$ CNN
$\bullet$ The model achieves best accuracy in fruits classifying $\bullet$ The fine-tuning process of the VGG-16 model requires significant computational resources and expertise to achieve optimal results. [213]
@@
34 Fruit Yield 2019 This paper presents the deep multi-faced system to increase banana fruits yield. $\bullet$ LSTM
$\bullet$ DL
$\bullet$ The model achieved better performance $\bullet$ Acquiring and curating such data can be time-consuming and resource-intensive [65]
@@
35 Fruit Detection 2020 FD should be robust to varying outdoor conditions with high signal to noise ratio. FD should help in estimating the angle at which the fruit is oriented around the stem for effective severing. $\bullet$ Semantic Segmentation using FCN
$\bullet$ Adaptive Image thresholding using video images using reinforced learning with decaying e-greedy algorithm
$\bullet$ FD system’s robustness and ability to handle varying outdoor conditions contribute to accurate fruit detection $\bullet$ Computational complexity [52]
@@
36 Fruit picking robots, 2020 The system possesses the ability to train with multiple fruit types, enabling accurate classification based on factors such as size and ripeness. $\bullet$ Improved efficiency $\bullet$ Multiple fruit crops support $\bullet$ Efficiency can be improved [53]
@@
37 Fruit picking 2020 This article presents a dual-arm aubergine harvesting robot consisting of two robotic arms configured in an anthropomorphic manner to optimize the dual workspace. $\bullet$ Cubic SVM
$\bullet$ water shed transformation
$\bullet$ Point cloud extraction
$\bullet$ The use of a dual-arm configuration in the aubergine harvesting robot allows for optimized workspace $\bullet$ Computational demands may limit real-time performance, requiring specialized hardware [55]
@@
38 Fruit Detectetion 2020 The system demonstrates effectiveness in detecting fruits even in environments where leaves or branches have similar colors, accommodating variability in fruit shape and irregularities in the growing environment. $\bullet$ CNN
$\bullet$ SVM
$\bullet$ Improved cucumber harvesting efficiency through effective fruit detection and high recognition accuracy $\bullet$ Limited adaptability for diverse agriculture applications [56]
39 Fruit Detectetion 2020 To locate & harvest cotton balls in field conditions. $\bullet$ Deep learning with Tiny Yolov3 with 7 conv. layers
$\bullet$ CLoDSA based image augmentation
@@
$\bullet$ Over $77\%$ Action Success Ratio (ASR) and over $97\%$ Detection Performance $\bullet$ Future design and development research should also include alternative energy sources to decrease energy costs [57]
@@
40 Fruit picking 2020 The paper introduces a harvesting robot designed specifically for plucking tea leaves, utilizing Jaco 2 robotic arm. $\bullet$ Jaco 2 based leaf plucking using ProMP $\bullet$ Increased efficiency and reduced labor in tea production. $\bullet$ Found limitations in handling branches with varying stiffness [59]
@@
41 Autonomous path detection 2020 The paper presents a system for autonomous path detection in semi-structured environments to enhance navigation and operation efficiency. $\bullet$ CNN $\bullet$ Potential for cost savings by minimizing human labor and errors. $\bullet$ Poor lighting, occlusions, and complex surroundings can limit the accuracy of autonomous path detection [60]
@@
42 Fruit Detection 2020 Improved detection of small harvesting cups using gray scale-depth and color images, with practical lighting-independent applications. $\bullet$ Faster RCNN with MobileNetV2 $\bullet$ Accurate detection of small harvesting cups in a wide range of lighting conditions, enabling more efficient and effective operations. $\bullet$ Dependency on color images for tapping position detection [169]
@@
43 Fruit Detection 2023 In this research, we propose a lightweight fruit detection algorithm for edge CPUs, enabling 19 FPS detection speed in RTFD-CPU smartphone app. $\bullet$ DNN
$\bullet$ PicoDet-S model
$\bullet$ The proposed RTFD has great potential for intelligent picking machines, $\bullet$ Sensitivity to environmental factors that may affect fruit detection accuracy [50]
@@
44 Better post harvest handling 2020 Improving harvest quality and productivity for farmers through innovative solutions. $\bullet$ Computer vision DL
$\bullet$ YOLOV3
$\bullet$ Assisting in the identification of healthy apples by distinguishing them from those with defects $\bullet$ Limitation in training different type of data [66]
@@
45 Better post harvest handling 2020 A ML approach utilizes radar back scattering data to accurately estimate grain quantity in silos. $\bullet$ KNN $\bullet$ Easy to apply in different grains $\bullet$ Challenges in accuaring correct radar data [68]
@@
46 Food security 2021 A hierarchical model is proposed, integrating ontology-enabled IoT, to distinguish and separate healthy Sekai-ichi apples using Boosted Continuous Non-spatial Whole Attribute Extraction (BCNAE). $\bullet$ IoT
$\bullet$ ML
$\bullet$ Achieving higher accuracy and throughput while minimizing processing time $\bullet$ Computational complexity and accuracy-speed trade-off are potential drawbacks when aiming for higher accuracy, throughput, and speed [78]
@@
47 Better post harvest handling 2022 The review explores the use of sensors, IoT principles, and ML technologies in post-harvest grain monitoring, identifying potential advancements and resource utilization. $\bullet$ IoT $\bullet$ Enhanced grain quality monitoring $\bullet$ Collecting sensor data requires robust measures to protect sensitive information. [67]
48 Food security 2022 The paper has investigated household-level food security. $\bullet$ RF
$\bullet$ decision tree
$\bullet$ Prone to over-fitting $\bullet$ The effectiveness of the model vary across different locations [75]
@@
49 Food security 2022 The paper has proposed detection and prediction models for identifying post-harvest deterioration zones in stored apple fruits. $\bullet$ CNN-based U-Net
$\bullet$ Deep-Lab
$\bullet$ Mask R-CNN
$\bullet$ The models enable retailers to maintain better quality control of their apple fruits, ensuring customer satisfaction. $\bullet$ The models are specifically designed for stored apple fruits only [76]
50 Food security 2022 The paper has presented Drone technology for food supply chain management. $\bullet$ DL
$\bullet$ ML
@@
$\bullet$ Reduces costs associated with traditional packing and delivery methods $\bullet$ Drones have limited payload capacity, battery life issues [77]
@@
Table 3 presents an overview of recent applications of ML techniques in the realm of agricultural practices, with a specific emphasis on genetic changes and plant breeding. The table chronicles research spanning from 2018 to 2023 and encapsulates various ML models. These models have been used to accelerate crop improvement, aid in precision decision making for rapid breeding, improve climate-resilient crop breeding, and even for automated plant phenotyping. Several references, such as [198] and [199], further emphasize the significance of ML in advancing modern agricultural practices and outcomes.
@@
Plant Breeding 2019 $\bullet$ Neural Network
$\bullet$ Explainable AI
@@
The author has presented the state of genomics and field phenomics and proposes a workable path to improvement. [198]
Plant Breeding 2023 $\bullet$ Ensemble Learning
$\bullet$ Big Data
$\bullet$ Random Forest
ML has been used in analyzing genetic data in plant breeding, accelerating crop improvement, developing new varieties, and ensuring food data security. [199]
Integrating Speed Breeding 2022 $\bullet$ Deep Learning
$\bullet$ Neural Network
@@
The author has utilized ML in speeding plant breeding and enabling precision decision-making. [200]
@@
Climate-Resilient Smart-Crop Breeding 2022 $\bullet$ Machine Learning The author has reviewed ML applications in crop breeding. [201]
Plant Breeding 2020 $\bullet$ Neural Network
$\bullet$ SVM
$\bullet$ Random Forest
Discussed improving new layers of plant breeding. [202]
Plant Breeding 2021 $\bullet$ Multilayer Perceptron
$\bullet$ SVM
$\bullet$ Random Forest
@@
ML models have been used for improving soybean crop yield. [203]
@@
Plant Breeding 2019 $\bullet$ Random Forest Predicted in-season seed yield (SY) in soybean based on high-dimensional phenotypic trait data. [204]
Plant Breeding 2019 $\bullet$ SVM
$\bullet$ Deep Learning
@@
Compared the performance of MLP and SVM methods to the Bayesian threshold genomic best linear unbiased prediction (TGBLUP) model for genomic selection in plant breeding. [205]
Estimating Phenotypic Traits 2018 $\bullet$ Regression
$\bullet$ Deep Learning
@@
Using color images and elevation maps, ML is employed to estimate emergence and biomass traits in wheat field plots. [193]
Phenotyping 2019 $\bullet$ Deep Learning
$\bullet$ CNN